### v2

@@ 5.1 QUBO Formulation
(1). $\textbf{b}_{j}$ is assigned to anchor $\textbf{a}_{i}$ by minimizing the function:
$f_{0}=\sum_{i=1}^{r}\sum_{j=1}^{s}\textbf{u}_{ij}d_{ij}$ (5)
where $d_{ij}$ is the cost between $\textbf{b}_{j}$ and $\textbf{a}_{i}$
(2). $\textbf{b}_{j}$ and $\textbf{b}_{k}$ can be assigned to the same anchor $\textbf{a}_{i}$ by minimizing the function:
$f_{1}=\sum_{i=1}^{r}\sum_{j=1}^{s}\sum_{k=1}^{s}\textbf{u}_{ij}d_{ijk}\textbf{u}_{ik}$ (6)
where $d_{ijk}$ is the triple costs among $\textbf{b}_{j}$, $\textbf{b}_{k}$ and $\textbf{a}_{i}$. It is measured by taking the average cost between each pair. That is $d_{ijk}=(d_{ij}+d_{ik}+d_{jk})/3$.
(3). We must minimize the following function to satisfy the constraint 4.
$f_{2}=\sum_{j=1}^{s}\sum_{i=1}^{r-1}\sum_{k=i+1}^{r}\textbf{u}_{ij}\textbf{u}_{kj}$ (7)
@@
$=\sum_{i=1}^{r}\sum_{j=1}^{s}\textbf{u}_{ij}d_{ij}+\sum_{i=1}^{r}\sum_{j=1}^{s}\sum_{k=1}^{s}\textbf{u}_{ij}d_{ijk}\textbf{u}_{ik}$
$+\alpha\sum_{j=1}^{s}\sum_{i=1}^{r-1}\sum_{k=i+1}^{r}\textbf{u}_{ij}\textbf{u}_{kj}$
The detailed algorithm of $\mathcal{P}$ is represented at the Algorithm 1.
Figure 2: An element of the similarity cost matrix is calculated by taking an averaged feature vector from the feature vectors in an anchor cluster and computing the negative cosine similarity with an unknown feature vector.
Algorithm 1 QUBO-based algorithm for clustering
  Input:     $\mathcal{G}=\{g_{0},g_{1},\dots,g_{u-1}\}$, processing batch: $b$
@@
  $\mathcal{A}\leftarrow\{\}$
@@
     $\mathcal{B}\leftarrow\textsf{rnd(}\mathcal{G},b$)
[…]
     $\mathcal{A}\leftarrow\textsf{Update(}\mathcal{A},\mathbf{u}\textsf{)}$
@@ 6.1 Experiment Settings
Datasets. This work is experimented on two datasets MNIST (Modified National Institute of Standards and Technology) [25] and CIFAR-100 (Canadian Institute for Advanced Research 100) [26]. MNIST is a database of handwritten digits commonly used as a benchmark dataset for image classification tasks. It contains a collection of 60,000 training images and 10,000 testing images. The images are grayscale and have a resolution of 28x28 pixels. Each image represents a handwritten digit from 0 to 9, labeled accordingly. CIFAR-100 is a more challenging dataset compared to MNIST. It comprises 60,000 color images, equally divided into 50,000 training and 10,000 testing images. The images have a resolution of 32x32 pixels and are divided into 100 different classes. Each class contains 500 training images and 100 testing images.
Deep Clustering Model. In our work, we have employed AlexNet [27] as our feature generator $\mathcal{M}$. AlexNet is a convolutional neural network architecture that gained significant attention and revolutionized the field of computer vision when it won the ImageNet Large Scale Visual Recognition Challenge in 2012. It consists of multiple layers, including convolutional layers, pooling layers, and fully connected layers. AlexNet is known for its deep architecture and rectified linear units (ReLU) as activation functions, which help capture complex features from input images. Using AlexNet as our feature generator, we leverage its ability to extract high-level and discriminative features from images. The feature dimension $d=64$. These features are then fed into the initial clustering algorithms $\mathcal{K}$. As defined in the previous works, we employ K-Nearish Neighbors (K-NN) algorithm as $\mathcal{K}$ to find $k=32$ closest neighbors of a datapoint and form it as a cluster. It is important to note that these clusters contain false positive samples. For this reason, we employ Clusformer [3] as a Deep Clustering Model to clean these clusters.
@@
$F_{B}=\frac{TP}{\sqrt{(TP+FP)\times(TP+FN)}}$ (9)
where $TP$ is the number of point pairs in the same cluster in both ground truth and prediction. $FP$ is the number of point pairs in the same cluster in ground truth but not in prediction. $FN$ is the number of point pairs in the same cluster in prediction but not in ground truth. Besides Pairwise F-score, BCubed-Fscore denoted as $F_{B}$ is also used for evaluation.
@@ 6.2 Experimental Results
Table 2: Comparison in face clustering performance on unlabeled images in CIFAR10 Database of our approach against prior approaches.
@@
To evaluate the effectiveness of the algorithms, we conducted clustering experiments on the CIFAR10 database and recorded the results in Table 1. The table provides a comprehensive overview of the clustering performance of K-Mean, HAC, and DBSCAN, allowing us to make meaningful comparisons. Interestingly, the Clusformer algorithm emerged as the standout performer, outperforming the previous methods by a substantial margin. Its superiority is evident across multiple performance metrics. For instance, Clusformer achieved an impressive 94.82% for $F_{P}$ (precision), 95.01% for $F_{B}$ (recall), and 92.38% for $NMI$, surpassing the results obtained by K-Mean, HAC, and DBSCAN by approximately 1.25% to 2.56%.
Moreover, we explored the application of Quantum Computing to enhance the performance of Clusformer. By leveraging the power of Quantum Computers to solve the Quadratic Unconstrained Binary Optimization (QUBO) problem, we achieved even more significant performance improvements. The utilization of QUBO resulted in Clusformer achieving remarkable scores of 96.20% for $F_{P}$, 96.75% for $F_{B}$, and 94.73% for $NMI$. These results solidify Clusformer’s superiority over previous methods and extend the improvement gap even further.
Similarly, We conduct the same experiment on the MNIST database as shown in Table 1. We obtained a similar conclusion as in the CIFAR10 database. It demonstrates that leveraging Quantum Machine to solve QUBO can further benefit the unsupervised clustering problem.
@@ 7 Conclusions
In this work, we have formulated a QUBO clustering algorithm for visual clustering and solved it efficiently using adiabatic quantum computing. We used experiments on the CIFAR-10 and MNIST datasets to demonstrate the performance of our QUBO algorithm. Considering current limitations inherent to the hardware of quantum computers, there is tremendous potential in the future to solve more complex and descriptive QUBO formulations using adiabatic quantum computing.

### v3

@@ 5.1 Cluster Constraint
In order to address this constraint, we propose the following optimization function
$f_{0}=\sum_{j=1}^{s}\sum_{i=1}^{r-1}\sum_{k=i+1}^{r}\textbf{u}_{ij}\textbf{u}_{kj}$ (6)
5.2 Cluster Assignment
The cluster $\textbf{b}_{j}$ is assigned to anchor $\textbf{a}_{i}$ if only if the cosine similarity distance, denoted as $d_{ij}$, is smallest. Since $\textbf{a}_{i}$ and $\textbf{b}_{j}$ are the cluster, we denote $\text{mean}(\textbf{a}_{i})$ is the center vector of the cluster $\textbf{a}_{i}$. Specifically, we measure the average of feature vectors in the cluster as the center. Therefore, the distance $d_{ij}$ can be formulated as $d_{ij}=\text{cosine}(\text{mean}(\textbf{b}_{j}),\text{mean}(\textbf{a}_{i}))$. For that reason, we define the following optimization function as in the Eqn (7):
$f_{1}=\sum_{i=1}^{r}\sum_{j=1}^{s}\textbf{u}_{ij}d_{ij}$ (7)
5.3 Triplet Assignment Constraint
In clustering, the assignment of data points to clusters often depends not just on individual similarities as in the Eqn (7) but also on pairwise relationships between data points. Let $\textbf{b}_{j}$ and $\textbf{b}_{k}$ are two separate clusters. They can be assigned to the same anchor $\textbf{a}_{i}$ by minimizing the function
$f_{2}=\sum_{i=1}^{r}\sum_{j=1}^{s}\sum_{k=1}^{s}\textbf{u}_{ij}d_{ijk}\textbf{u}_{ik}$ (8)
where $d_{ijk}$ is the triple costs among $\textbf{b}_{j}$, $\textbf{b}_{k}$ and $\textbf{a}_{i}$. It is measured by taking the average cost between each pair. That is $d_{ijk}=(d_{ij}+d_{ik}+d_{jk})/3$. This equation accounts for the pairwise dependencies, ensuring that similar data points are grouped in the same cluster. In addition, $f_{1}$ is inherently quadratic due to the product of binary variables $\textbf{u}_{ij}$ and $\textbf{u}_{jk}$, which is key for encoding problems into QUBO. It makes $f_{1}$ directly suitable for optimization using quantum annealers or other solvers designed for QUBO problems. By minimizing $f_{1}$, the model encourages data points with a strong relationship (low $d_{ijk}$) to be assigned to the same cluster anchor $\textbf{a}_{i}$. It enhances intra-cluster similarity and improves clustering quality.
5.4 Overall Optimization Function
@@
$=\sum_{j=1}^{s}\sum_{i=1}^{r-1}\sum_{k=i+1}^{r}\textbf{u}_{ij}\textbf{u}_{kj}$
$+\sum_{i=1}^{r}\sum_{j=1}^{s}\textbf{u}_{ij}d_{ij}+\sum_{i=1}^{r}\sum_{j=1}^{s}\sum_{k=1}^{s}\textbf{u}_{ij}d_{ijk}\textbf{u}_{ik}$ (10)
5.5 Pseudo Algorithm
We have formulated the clustering component $\mathcal{P}$ as in the QUBO form. The details process of $\mathcal{P}$ is represented in the Algorithm 1. In particular, the algorithm demonstrates steps to construct the anchor clusters $\mathcal{A}$ from a set of clusters $\mathcal{G}$. Firstly, we random sampling $b$ clusters from $\mathcal{G}$. Next, we solve the QUBO to find the assignments. Finally, we extend the anchor cluster $\textbf{a}\in\mathcal{A}$ by the new assignments.
Algorithm 1 QUBO-based Clustering Algorithm
  Input:$\mathcal{G}=\{g_{0},g_{1},\dots,g_{u-1}\}$: Set of elements to cluster  $b$: Batch size for processing
@@ 5.5 Pseudo Algorithm
  Initialize an empty set of anchor clusters: $\mathcal{A}\leftarrow\{\}$
@@
     Select a random subset (batch) of elements: $\mathcal{B}\leftarrow\textsf{SelectBatch}(\mathcal{G},b)$
     Solve the QUBO problem for batch clustering: $\mathbf{u}\leftarrow\textsf{SolveQUBO}(\mathcal{A},\mathcal{B})$
     Update anchor clusters with the new assignments: $\mathcal{A}\leftarrow\textsf{UpdateClusters}(\mathcal{A},\mathbf{u})$
     Remove processed batch $\mathcal{B}$ from $\mathcal{G}$: $\mathcal{G}\leftarrow\textsf{RemoveBatch}(\mathcal{G},\mathcal{B})$
@@ 6.1 Experiment Settings
Datasets. This work is experimented on two datasets MNIST (Modified National Institute of Standards and Technology) [25] and CIFAR100 [26]. MNIST is a database of handwritten digits commonly used as a benchmark dataset for image classification tasks. It contains a collection of 60,000 training images and 10,000 testing images. The images are grayscale and have a resolution of 28x28 pixels. Each image represents a handwritten digit from 0 to 9, labeled accordingly. CIFAR100 is a more challenging dataset compared to MNIST. It comprises 60,000 color images, equally divided into 50,000 training and 10,000 testing images. The images have a resolution of $32\times 32$ pixels and are divided into 100 different classes. Each class contains 500 training images and 100 testing images.
Deep Clustering Model. In our work, we have employed AlexNet [27] as our feature generator $\mathcal{M}$. AlexNet is a convolutional neural network architecture that gained significant attention and revolutionized the field of computer vision when it won the ImageNet Large Scale Visual Recognition Challenge in 2012. It consists of multiple layers, including convolutional, pooling, and fully connected layers. AlexNet is known for its deep architecture and rectified linear units (ReLU) as activation functions, which help capture complex features from input images. Using AlexNet as our feature generator, we leverage its ability to extract high-level and discriminative features from images. The feature dimension $d=64$. These features are then fed into the initial clustering algorithms $\mathcal{K}$. As defined in the previous works, we employ the K-Nearish Neighbors (K-NN) algorithm as $\mathcal{K}$ to find $k=32$ closest neighbors of a datapoint and form it as a cluster. It is important to note that these clusters contain false positive samples. For this reason, we employ Clusformer [3] (as the role of $\mathcal{M}$ in Eqn (2)) which is a transformer-based architecture designed for unsupervised large-scale recognition such as facial or visual landmark clustering. The goal of Clusformer is to detect noisy or hard samples inside clusters. This approach promises to refine the clustering results from K-NN.
@@
$F_{P}=\frac{TP}{\sqrt{(TP+FP)\times(TP+FN)}}$ (11)
where $TP$ is the number of point pairs in the same cluster in both ground truth and prediction. $FP$ is the number of point pairs in the same cluster in ground truth but not prediction. $FN$ is the number of point pairs in the same cluster in prediction but not in ground truth. Besides Pairwise F-score, BCubed-Fscore denoted as $F_{B}$ is also used for evaluation.
@@ 6.2 Experimental Results
Table 2: Comparison in face clustering performance on unlabeled images in CIFAR100 Database of our approach against prior approaches.
@@
To evaluate the effectiveness of the algorithms, we conducted clustering experiments on the CIFAR100 database and recorded the results in Table 1. The table provides a comprehensive overview of the clustering performance of K-Mean, HAC, and DBSCAN, allowing us to make meaningful comparisons. Interestingly, the Clusformer algorithm emerged as the standout performer, outperforming the previous methods by a substantial margin. Its superiority is evident across multiple performance metrics. For instance, Clusformer achieved an impressive 94.82% for $F_{P}$ (precision), 95.01% for $F_{B}$ (recall), and 92.38% for $NMI$, surpassing the results obtained by K-Mean, HAC, and DBSCAN by approximately 1.25% to 2.56%.
Moreover, we explored the application of Quantum Computing to enhance the performance of Clusformer. By leveraging the power of Quantum Computers to solve the Quadratic Unconstrained Binary Optimization (QUBO) problem, we achieved even more significant performance improvements. The utilization of QUBO resulted in Clusformer achieving remarkable scores of 96.20% for $F_{P}$, 96.75% for $F_{B}$, and 94.73% for $NMI$. These results solidify Clusformer’s superiority over previous methods and further extend the improvement gap.
Similarly, We conduct the same experiment on the CIFAR database as shown in Table 2. We obtained a similar conclusion as in the CIFAR100 database. It demonstrates that leveraging Quantum Machine to solve QUBO can further benefit the unsupervised clustering problem.
@@ 7 Conclusions
In this work, we have formulated a QUBO clustering algorithm for visual clustering and solved it efficiently using adiabatic quantum computing. We used experiments on the CIFAR100 and MNIST datasets to demonstrate the performance of our QUBO algorithm. Considering current limitations inherent to the hardware of quantum computers, there is tremendous potential in the future to solve more complex and descriptive QUBO formulations using adiabatic quantum computing.
8 Availability of data and materials
The data that supports the findings of this study can be found at https://www.cs.toronto.edu/ kriz/cifar.html and http://yann.lecun.com/exdb/mnist/
Author Biography
Xuan-Bac Nguyen is currently a Ph.D. student at the Department of Computer Science and Computer Engineering of the University of Arkansas. He received his M.Sc. degree in Computer Science from the Electrical and Computer Engineering Department at Chonnam National University, South Korea, in 2020. He received his B.Sc. degree in Electronics and Telecommunications from the University of Engineering and Technology, VNU, in 2015. In 2016, he was a software engineer in Yokohama, Japan. His research interests include Quantum Machine Learning, Face Recognition, Facial Expression, and Medical Image Processing.
Hugh Churchill is a Professor and 21st Century Chair in Nanophysics in the Department of Physics at the University of Arkansas (UA). He is also Associate Director for Operations of the MonArk NSF Quantum Foundry. He received a Ph.D. in Physics from Harvard University in 2012 and was a Pappalardo postdoctoral fellow in Physics at MIT. Since 2015 he has led the Quantum Device Laboratory at UA, pursuing research interests in quantum materials and devices, low-dimensional materials, quantum transport, optoelectronics, and automation of experiments.
Khoa Luu is an Assistant Professor and the Director of the Computer Vision and Image Understanding (CVIU) Lab in the Department of Electrical Engineering and Computer Science (EECS) at the University of Arkansas (UA), Fayetteville, US. He is affiliated with the NSF MonARK Quantum Foundry. He is an Associate Editor of the IEEE Access Journal and the Multimedia Tools and Applications Journal, Springer Nature. He is also the Area Chair in CVPR 2023, CVPR 2024, NeurIPS 2024, WACV 2025, and ICLR 2025. He was the Research Project Director at the Cylab Biometrics Center at Carnegie Mellon University (CMU), USA. His research interests focus on various topics, including Quantum Machine Learning, Biometrics, Smart Health, and Precision Agriculture. He has received eight patents and three Best Paper Awards and coauthored 120+ papers in conferences, technical reports, and journals. He was a Vice-Chair of the Montreal Chapter of the IEEE Systems, Man, and Cybernetics Society in Canada from September 2009 to March 2011. He was the Technical Program Chair at the IEEE GreenTech Conference 2024 and a co-organizer and chair of the CVPR Precognition Workshop in 2019-2024, MICCAI Workshop in 2019 and 2020, and ICCV Workshop in 2021. He was a PC member of AAAI, ICPRAI in 2020, 2021, 2022. He has been an active reviewer for several top-tier conferences and journals, such as CVPR, ICCV, ECCV, NeurIPS, ICLR, FG, BTAS, IEEE-TPAMI, IEEE-TIP, IEEE Access, Journal of Pattern Recognition, Journal of Image and Vision Computing, Journal of Signal Processing, and Journal of Intelligence Review.
Samee U. Khan received a Ph.D. in 2007 from the University of Texas, Arlington, TX. Currently, he is Professor of Electrical & Computer Engineering at Mississippi State University (MSU) and served as Department Head from Aug. 2020 to July 2024. Before arriving at MSU, he was Cluster Lead (2016-2020) for Computer Systems Research at the National Science Foundation and the Walter B. Booth Professor at North Dakota State University. His research interests include optimization, robustness, and security of computer systems. His work has appeared in over 450 publications. He is the associate editor of IEEE Transactions on Cloud Computing and the Journal of Parallel and Distributed Computing. He is a Fellow of the IET and BCS. He is a Distinguished Member of the ACM and a Senior Member of the IEEE. He has won several awards, including the Best Paper Award (Systems Track), IEEE Cloud Summit, 2024; IEEE R3 Outstanding Engineer Award, 2024; IEEE Computer Society Distinguished Contributor Award, 2022 (inducted in the inaugural class); IEEE ComSoc Technical Committee on Big Data Best Journal Paper Award, 2019; IEEE-USA Professional Achievement Award, 2016; IEEE Golden Core Member Award, 2016; IEEE TCSC Award for Excellence in Scalable Computing Research (Middle Career Researcher), 2016; IEEE Computer Society Meritorious Service Certificate, 2016; Tapestry of Diverse Talents Award, North Dakota State University (NDSU), ND, USA, 2016; Exemplary Editor, IEEE Communications Surveys and Tutorials, IEEE Communications Society, 2014; Outstanding Summer Undergraduate Research Faculty Mentor Award, NDSU, ND, USA, 2013; Best Paper Award, IEEE Intl. Conf. on Scalable Computing and Communications (ScalCom), 2012; Sudhir Mehta Memorial International Faculty Award, NDSU, ND, USA, 2012; Best Paper Award, ACM/IEEE Intl. Conf. on Green Computing & Communications (GreenCom), 2010.